# Markov switching zero-inflated space-time multinomial models for comparing multiple infectious diseases

### v1

@@ 2.1 Incorporating zero-inflation
As detailed in the introduction, if some of the diseases are absent in an area then the ARMN model is not appropriate as it will always assign a positive probability to cases being reported. In this subsection, we will keep the definitions from above and extend the ARMN model to deal with long periods of disease absence by adapting the proposal of Zeng et al. (2022) to our space-time setting. Let $S_{kit}$, for $k=1,\dots,K$, be an indicator for the presence of disease $k$, so that $S_{kit}=1$ if disease $k$ was present in area $i$ during time $t$ and $S_{kit}=0$ if disease $k$ was absent. Under the zero-inflated multinomial framework of Zeng et al. (2022), the multinomial probabilities are given by,
@@
for $k=2,\dots,K$. An immediate issue with (5) is that we cannot have $S_{1it}=\dots=S_{Kit}=0$ as we would divide by 0. Even if we were to define $\bm{\pi}_{it}=\bm{0}$ if $S_{1it}=\dots=S_{Kit}=0$, we would have the issue that this implies $\text{total}_{it}=0$. A multinomial model conditions on the total count and does not generate it. There is nothing in the model of Zeng et al. (2022) to prevent $S_{1it}=\dots=S_{Kit}=0$. They may have overlooked this as they dealt with the case of $K\geq 50$ and, therefore, it is very unlikely $S_{1it}=\dots=S_{Kit}=0$. However, in our motivating example, we have $K=3$.
To prevent $S_{1it}=\dots=S_{Kit}=0$, as it makes sense in many epidemiological applications, we will assume that the baseline disease is well-established in the region and is always present. That is, we assume $S_{1it}=1$ for all $i$ and $t$. For instance, in our motivating example, dengue has been circulating in Rio since 1986 and is well-established there (Teixeira et al., 2009). We replace Equation (5) then by,
$\begin{split}\pi_{1it}=\frac{1}{1+\sum_{j=2}^{K}S_{jit}\lambda_{jit}^{*}},\hskip 28.45274pt\pi_{kit}=\frac{S_{kit}\lambda_{kit}^{*}}{1+\sum_{j=2}^{K}S_{jit}\lambda_{jit}^{*}},\end{split}$ (6)
for $k=2,\dots,K$.
The presence of a disease in an area should depend on whether it was present previously, as disease presence and absence are often highly persistent (Coutinho et al., 2006; Douwes-Schultz and Schmidt, 2022), and may also depend on whether other diseases were present previously, as the diseases can interact with one another (Paul et al., 2008; Sherlock et al., 2013). Also, the presence of a disease might depend on environmental or socioeconomic factors, like temperature and water supply in the case of vector-borne diseases (Schmidt et al., 2011; Xu et al., 2017). Therefore, we model $S_{kit}$ as,
@@
where $\bm{z}_{kit}$ is a vector of space-time covariates that might affect the presence of disease $k$. As the disease may spread from neighboring areas (Grenfell et al., 2001), we allow functions of disease cases in neighboring areas to be added to $\bm{z}_{kit}$, hence the dependence on $\bm{y}_{t-1}$ in (7). For instance, the prevalence of the disease in neighboring areas could be added, see Section 4. Note, as $\rho_{k}^{AR}$ in Equation (7) increases and $\eta_{0k}$ decreases, the probability of getting a consecutive period of disease presence or absence approaches 1. Therefore, (7) can account for long consecutive periods of disease presence and absence which are often observed for infectious diseases, see Figure 1 and Coutinho et al. (2006).
We will call the ARMN model modified by Equations (6)-(7) the Markov switching zero-inflated autoregressive multinomial (MS-ZIARMN) model. Note, from Equation (6), $\pi_{1it}+\dots+\pi_{Kit}=1$ for any possible values of the parameters and presence indicators. Also note that if diseases $k$ and $j$ are present in area $i$ during time $t$ we have $\pi_{kit}/\pi_{jit}=\lambda_{kit}^{*}/\lambda_{jit}^{*}$, which is the same as in the ARMN model described in Equations (2)-(4). That is, the MS-ZIARMN model preserves the relative distribution of the present diseases from the ARMN model. Finally, if $K=2$ and we assume Equation (6), where $\log(\lambda_{2it}^{*})=\text{logit}(\pi_{2it})=\alpha_{02i}+\bm{x}_{2it}^{T}\bm{\alpha}_{2}$ and $\text{logit}(p_{2it})=\eta_{02}+\bm{z}_{2it}^{T}\bm{\eta}_{2}$, we get the well known and popular zero-inflated binomial (ZIB) model of Hall (2000). Therefore, our model without space-time effects could be seen as a multinomial extension of Hall (2000).
@@ 3 Inferential Procedure
Here we assume the number of diseases $K$ is small enough for matrix multiplication with a $2^{K}$ by $2^{K}$ matrix to be computationally feasible. This should be the case in most epidemiological applications, for instance, in our motivating example we compare $K=3$ diseases. Our inferential strategy revolves around reparametrizing the MS-ZIARMN model as a Markov switching model (Frühwirth-Schnatter, 2006), which allows for very efficient Bayesian inference using the forward filtering backward sampling (FFBS) algorithm (Chib, 1996). To simplify the explanations we will assume, without loss of generality, that $K=3$.
In the case of $K=3$, we have $\bm{S}_{it}=(S_{2it},S_{3it})^{T}$ as $S_{1it}$ is always equal to 1. Let $S_{it}^{*}$ be an indicator for the possible values of $\bm{S}_{it}$, so that $S_{it}^{*}=1$ if $\bm{S}_{it}=(1,1)^{T}$, $S_{it}^{*}=2$ if $\bm{S}_{it}=(0,1)^{T}$, $S_{it}^{*}=3$ if $\bm{S}_{it}=(1,0)^{T}$ and $S_{it}^{*}=4$ if $\bm{S}_{it}=(0,0)^{T}$. Then, from Equation (6),
$\bm{y}_{it}|\bm{y}_{t-1},S_{it}^{*},\text{total}_{it}\sim\begin{cases}\text{Multinom}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1}{1+\lambda_{2it}^{*}+\lambda_{3it}^{*}},\frac{\lambda_{2it}^{*}}{1+\lambda_{2it}^{*}+\lambda_{3it}^{*}},\frac{\lambda_{3it}^{*}}{1+\lambda_{2it}^{*}+\lambda_{3it}^{*}}}\right)^{T},\text{total}_{it}}\right),&\text{ if $S_{it}^{*}=1$}\\[5.0pt] \text{Multinom}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1}{1+\lambda_{3it}^{*}},0,\frac{\lambda_{3it}^{*}}{1+\lambda_{3it}^{*}}}\right)^{T},\text{total}_{it}}\right),&\text{ if $S_{it}^{*}=2$}\\[5.0pt] \text{Multinom}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1}{1+\lambda_{2it}^{*}},\frac{\lambda_{2it}^{*}}{1+\lambda_{2it}^{*}},0}\right)^{T},\text{total}_{it}}\right),&\text{ if $S_{it}^{*}=3$}\\[5.0pt] \text{Multinom}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(1,0,0}\right)^{T},\text{total}_{it}}\right),&\text{ if $S_{it}^{*}=4$},\end{cases}$ (8)
@@
for $\bm{s}_{i(t-1)},\bm{s}_{it}=(1,1),(0,1),(1,0),(0,0)$, using Equation (7). Finally, as $S_{i2}^{*}$ depends on $S_{i1}^{*}$ we require an initial distribution for the Markov chain, that is, $P(S_{i1}^{*}=j)$ for $j=1,2,3,4$ and $i=1,\dots,N$. The modeler can set the initial distribution based on how likely they believe the diseases to be present at the beginning of the study period.
Given $p(\bm{y}_{it}|\bm{y}_{t-1},S_{it}^{*},\text{total}_{it})$ and $\Gamma(S_{it}^{*}|\bm{y}_{t-1})$ for $i=1,\dots N$ and $t=2,\dots T$, and $p(S_{i1}^{*})$ for $i=1,\dots,N$, we completely define a Markov switching model (Frühwirth-Schnatter, 2006). A Markov switching model assumes a time series can be described by several submodels, often called states or regimes, where switching between submodels is governed by a first-order Markov chain (Hamilton, 1989). In our case, the submodels are given by the 4 multinomial distributions in (8) and we switch between these submodels through the transition matrix $\Gamma(S_{it}^{*}|\bm{y}_{t-1})$. Typically, $S_{it}^{*}$ is called the state indicator.
Let $\bm{S}^{*}=(\bm{S}_{1}^{*},\dots,\bm{S}_{T}^{*})^{T}$ be the vector of all state indicators, where $\bm{S}_{t}^{*}=(S_{1t}^{*},\dots,S_{Nt}^{*})^{T}$, $\bm{y}=(\bm{y}_{1},\dots,\bm{y}_{T})^{T}$ be the vector of all observations, and $\bm{\beta}$ the vector of all parameters in the multinomial part of the model (see (2) and (4)); further, let $\bm{\theta}$ be the vector of all parameters in the Markov chain part of the model (see (7)), and, finally, let $\bm{v}=(\bm{\beta},\bm{\theta})^{T}$ be the vector of all model parameters. The likelihood of $\bm{v}$ given $\bm{S}^{*}$ and $\bm{y}$ is given by,
@@
Note, from Equation (8), $S_{it}^{*}$ is only fully known when $y_{1it},y_{2it},y_{3it}>0$, in which case we must have $S_{it}^{*}=1$. It is possible, assuming $K$ is not too large, to use the forward filter (Hamilton, 1989) to completely marginalize $\bm{S}^{*}$ from (9) and calculate $p(\bm{y}|\bm{v})$, see SM Section 3. However, we want to make inferences about $\bm{S}^{*}$ to investigate when the model believes the diseases were present or absent for model checking, see SM Section 5. Therefore, $\bm{S}^{*}$ is estimated along with $\bm{v}$ by sampling both from their joint posterior distribution $p(\bm{v},\bm{S}^{*}|\bm{y})\propto p(\bm{S}^{*},\bm{y}|\bm{v})p(\bm{v})$, where $p(\bm{v})$ is the prior distribution of $\bm{v}$. We specified wide independent priors for most lower-level elements of $\bm{v}$. We specified a conjugate inverse-Wishart prior for $\bm{\Sigma}$ with $K$ degrees of freedom and identity matrix, as it implies the correlations in $\bm{\Sigma}$ have marginal uniform prior distributions (Gelman et al., 2013).
@@
As the joint posterior is not available in closed form, we resorted to Markov chain Monte Carlo (MCMC) methods, in particular, we used a hybrid Gibbs sampling algorithm with some steps of the Metropolis-Hastings algorithm to sample from it. The full details of the Gibbs sampler are given in SM Section 3. We sampled all of $\bm{S}^{*}$ jointly from $p(\bm{S}^{*}|\bm{v},\bm{y})$ using the FFBS algorithm for Markov switching models (Chib, 1996). We found our MCMC algorithm mixed much faster when jointly sampling $\bm{S}^{*}$ compared to sampling each presence indicator individually. This is typically the case for Markov switching models (Frühwirth-Schnatter, 2006) and was one of our main motivations for reparametrizing the MS-ZIARMN model as a Markov switching model.
@@
Our MCMC algorithm was implemented using the R package Nimble (de Valpine et al., 2017). We implemented the FFBS algorithm using Nimble’s custom sampler feature. All other MCMC samplers mentioned in SM Section 3 are built into Nimble. All code and data are available on GitHub https://github.com/Dirk-Douwes-Schultz/MS_ZIARMN_code.
4 Application to Counts of Dengue, Zika and Chik.
@@
Between 2015 and 2016 Rio de Janeiro experienced its first triple epidemic of dengue, Zika and chikungunya. Using space-time cluster analysis, Freitas et al. (2019) found that Zika and dengue clusters were much more likely in the western region of Rio (there were no chikungunya clusters found in the west), while other regions contained many clusters of each disease though they often differed in time. However, they did not investigate how these differences in clustering might depend on covariates or other factors. Schmidt et al. (2022), using a spatial multinomial regression model, found that cases of chikungunya were more likely in urban areas and Zika cases were less likely in population-dense areas, compared to dengue cases. A spatial analysis cannot investigate how space-time factors might have affected differences in the transmission intensity of the diseases, which may be important in this example, see below. Therefore, we apply our MS-ZIARMN model to cases of the three diseases, $K=3$, across the 160 neighborhoods of Rio de Janeiro, $N=160$, for each bi-week between 2015-2016, $T=52$. Our model seems appropriate for this example as there appear to be long periods of disease absence in both the time series of chikungunya and Zika cases, see Figure 1 for example (overall 65% of Zika cases are equal to 0 and 75% of chikungunya cases are equal to 0).

### v2

@@ 2.1 Incorporating zero-inflation
Let $S_{kit}$, for $k=1,\dots,K$, be an indicator for the presence of disease $k$, so that $S_{kit}=1$ if disease $k$ was present in area $i$ during time $t$ and $S_{kit}=0$ if disease $k$ was absent. Following traditional zero-inflated epidemiological models (Vergne et al., 2014; Douwes-Schultz and Schmidt, 2022), we will assume that if a disease is absent ($S_{kit}=0$), no cases will be reported (a true or structural 0). Note that this means if all diseases are absent at once in an area ($S_{1it}=S_{2it}\dots=S_{Kit}=0$), the total number of cases would be 0 ($\text{total}_{it}=0$). However, a multinomial model conditions on the total count and does not generate it. Therefore, as it makes sense in many epidemiological applications, we will assume that the baseline disease is well established in the region and always present. That is, we assume $S_{1it}=1$ for all $i$ and $t$. We then assume that if a disease is present, the odds of a new case being from that disease relative to the baseline disease is given by Equation (2) above, leading to the following zero-inflated multinomial model,
[…]
where $\bm{S}_{it}=(S_{2it},\dots,S_{Kit})^{T}$. Note from (5) that we can have zero-reported cases of a disease ($y_{kit}=0$) even if the disease is present ($S_{kit}=1$). Following Vergne et al. (2014) and Douwes-Schultz and Schmidt (2022), we interpret this as the disease being present but undetected by the surveillance system (a false or at-risk 0).
We can compare (5) with the zero-inflated multinomial framework of Zeng et al. (2022), who assumed
@@
for $k=2,\dots,K$. An immediate issue with (6) is that we cannot have $S_{1it}=\dots=S_{Kit}=0$ as we would divide by 0. Even if we were to define $\bm{\pi}_{it}=\bm{0}$ if $S_{1it}=\dots=S_{Kit}=0$, we would have the same issue that this implies $\text{total}_{it}=0$. There is nothing in the model of Zeng et al. (2022) to prevent $S_{1it}=\dots=S_{Kit}=0$. They may have overlooked this as they dealt with the case of $K\geq 50$ and, therefore, it is very unlikely $S_{1it}=\dots=S_{Kit}=0$. However, in our motivating example, we have $K=3$. Note that Zeng’s model (6) assuming $S_{1it}=1$ for all $i$ and $t$ is equivalent to (5).
The presence of a disease in an area should depend on whether it was present previously, as disease presence and absence are often highly persistent (Coutinho et al., 2006; Douwes-Schultz and Schmidt, 2022), and may also depend on whether other diseases were present previously, as the diseases can interact with one another (Paul et al., 2008; Sherlock et al., 2013). Additionally, the presence of a disease may depend on environmental or socioeconomic factors, such as temperature and water supply in the case of vector-borne diseases (Schmidt et al., 2011; Xu et al., 2017). Therefore, we model $S_{kit}$ as,
@@
where $\bm{z}_{kit}$ is a vector of space-time covariates that might affect the presence of disease $k$. As the disease may spread from neighboring areas (Grenfell et al., 2001), we allow functions of disease cases in neighboring areas to be added to $\bm{z}_{kit}$, hence the dependence on $\bm{y}_{t-1}$ in (7). For instance, the prevalence of the disease in neighboring areas could be added, see Section 4. Note, as $\rho_{k}^{AR}$ in Equation (7) increases and $\eta_{0k}$ decreases, the probability of getting a consecutive period of disease presence or absence approaches 1. Therefore, (7) can account for long consecutive periods of disease presence and absence, which are often observed for infectious diseases (see Figure 1 and Coutinho et al. (2006)).
We will call the ARMN model modified by Equations (5) and (7) the Markov switching zero-inflated autoregressive multinomial (MS-ZIARMN) model. Note, from Equation (5), $\pi_{1it}+\dots+\pi_{Kit}=1$ for any possible values of the parameters and presence indicators. Also note that if diseases $k$ and $j$ are present in area $i$ during time $t$ we have $\pi_{kit}/\pi_{jit}=\lambda_{kit}^{*}/\lambda_{jit}^{*}$, which is the same as in the ARMN model described in Equations (2)-(4). That is, the MS-ZIARMN model preserves the relative distribution of the present diseases from the ARMN model. Finally, if $K=2$ and we assume Equation (5), where $\log(\lambda_{2it}^{*})=\text{logit}(\pi_{2it})=\alpha_{02i}+\bm{x}_{2it}^{T}\bm{\alpha}_{2}$ and $\text{logit}(p_{2it})=\eta_{02}+\bm{z}_{2it}^{T}\bm{\eta}_{2}$, we get the well known and popular zero-inflated binomial (ZIB) model of Hall (2000). Therefore, our model without space-time effects could be seen as a multinomial extension of Hall (2000).
@@ 3 Inferential Procedure
Here, we assume the number of diseases $K$ is small enough for matrix multiplication with a $2^{K}$ by $2^{K}$ matrix to be computationally feasible. This should be the case in most epidemiological applications; for instance, in our motivating example, we compare $K=3$ diseases. Our inferential strategy revolves around reparametrizing the MS-ZIARMN model as a Markov switching model (Frühwirth-Schnatter, 2006), which allows for very efficient Bayesian inference using the forward filtering backward sampling (FFBS) algorithm (Chib, 1996). To simplify the explanations, we will assume, without loss of generality, that $K=3$.
In the case of $K=3$, we have $\bm{S}_{it}=(S_{2it},S_{3it})^{T}$. Let $S_{it}^{*}$ be an indicator for the possible values of $\bm{S}_{it}$, so that $S_{it}^{*}=1$ if $\bm{S}_{it}=(1,1)^{T}$, $S_{it}^{*}=2$ if $\bm{S}_{it}=(0,1)^{T}$, $S_{it}^{*}=3$ if $\bm{S}_{it}=(1,0)^{T}$ and $S_{it}^{*}=4$ if $\bm{S}_{it}=(0,0)^{T}$. Then, from Equation (5),
[…]
@@
for $\bm{s}_{i(t-1)},\bm{s}_{it}=(1,1),(0,1),(1,0),(0,0)$, using Equation (7). Finally, as $S_{i2}^{*}$ depends on $S_{i1}^{*}$, we require an initial distribution for the Markov chain, that is, $P(S_{i1}^{*}=j)$ for $j=1,2,3,4$ and $i=1,\dots,N$. The modeler can set the initial distribution based on how likely they believe the diseases to be present at the beginning of the study period.
Given $p(\bm{y}_{it}|\bm{y}_{t-1},S_{it}^{*},\text{total}_{it})$ and $\Gamma(S_{it}^{*}|\bm{y}_{t-1})$ for $i=1,\dots N$ and $t=2,\dots T$, and $p(S_{i1}^{*})$ for $i=1,\dots,N$, we completely define a Markov switching model (Frühwirth-Schnatter, 2006). A Markov switching model assumes a time series can be described by several submodels, often called states or regimes, where switching between submodels is governed by a first-order Markov chain (Hamilton, 1989). In our case, the submodels are given by the 4 multinomial distributions in (8), and we switch between these submodels through the transition matrix $\Gamma(S_{it}^{*}|\bm{y}_{t-1})$. Typically, $S_{it}^{*}$ is called the state indicator.
Let $\bm{S}^{*}=(\bm{S}_{1}^{*},\dots,\bm{S}_{T}^{*})^{T}$ be the vector of all state indicators, where $\bm{S}_{t}^{*}=(S_{1t}^{*},\dots,S_{Nt}^{*})^{T}$; $\bm{y}=(\bm{y}_{1},\dots,\bm{y}_{T})^{T}$ be the vector of all observations; and $\bm{\beta}$ the vector of all parameters in the multinomial part of the model (see (2)-(4)). Further, let $\bm{\theta}$ be the vector of all parameters in the Markov chain part of the model (see (7)), and, finally, let $\bm{v}=(\bm{\beta},\bm{\theta})^{T}$ be the vector of all model parameters. The joint distribution of $\bm{S}^{*}$ and $\bm{y}$ given $\bm{v}$ is given by,
@@
Note, from Equation (8), $S_{it}^{*}$ is only fully known when $y_{1it},y_{2it},y_{3it}>0$, in which case we must have $S_{it}^{*}=1$. It is possible, assuming $K$ is not too large, to use the forward filter (Hamilton, 1989) to completely marginalize $\bm{S}^{*}$ from (9) and calculate $p(\bm{y}|\bm{v})$, which as a function of $\bm{v}$ is the marginal likelihood function, see SM Section 3. However, we want to make inferences about $\bm{S}^{*}$ to investigate when the model believes the diseases were present or absent for model checking, as detailed in SM Section 6. Therefore, $\bm{S}^{*}$ is estimated along with $\bm{v}$ by sampling both from their joint posterior distribution $p(\bm{v},\bm{S}^{*}|\bm{y})\propto p(\bm{S}^{*},\bm{y}|\bm{v})p(\bm{v})$, where $p(\bm{v})$ is the prior distribution of $\bm{v}$. We will now discuss the prior specification in detail.
We specified a conjugate inverse-Wishart prior for $\bm{\Sigma}$ with $K$ degrees of freedom and identity matrix, as it implies the correlations in $\bm{\Sigma}$ have marginal uniform prior distributions (Gelman et al., 2013). For all non-bounded parameters, such as the intercepts and covariate effects, we used wide normal priors with a standard deviation of 5. For the precisions $1/\sigma^{2}_{k}$ we used conjugate $\text{gamma}(.1,.1)$ priors.
@@
As the joint posterior is not available in closed form, we resorted to Markov chain Monte Carlo (MCMC) methods; in particular, we used a hybrid Gibbs sampling algorithm with some steps of the Metropolis-Hastings algorithm to sample from it. The full details of the Gibbs sampler are given in SM Section 3. We sampled all of $\bm{S}^{*}$ jointly from $p(\bm{S}^{*}|\bm{v},\bm{y})$ using the FFBS algorithm for Markov switching models (Chib, 1996). We found our MCMC algorithm mixed much faster when jointly sampling $\bm{S}^{*}$ compared to sampling each presence indicator individually. This is typically the case for Markov switching models (Frühwirth-Schnatter, 2006), and was one of our main motivations for reparametrizing the MS-ZIARMN model as a Markov switching model.
@@
Our MCMC algorithm was implemented using the R package Nimble (de Valpine et al., 2017). We implemented the FFBS algorithm using Nimble’s custom sampler feature. All other MCMC samplers mentioned in SM Section 3 are built into Nimble. All code and data are available on GitHub https://github.com/Dirk-Douwes-Schultz/MS_ZIARMN_code. In SM Section 4, we provide a simulation study, which shows that our proposed Gibbs sampler can recover the true parameters of an MS-ZIARMN model that is specified like in our motivating example in Section 4.
4 Application to Counts of Dengue, Zika and Chikungunya
@@
Since 2015, several countries in Latin America have experienced triple epidemics of dengue, Zika and chikungunya, placing an immense burden on the local populations (Rodriguez-Morales et al., 2016; Bisanzio et al., 2018). Between 2015 and 2016, Rio de Janeiro experienced its first triple epidemic of these diseases. Despite being transmitted by the same vector, the Aedes aegypti, Freitas et al. (2019), using space-time cluster detection methods, found that Zika and dengue clusters were much more likely in the western region of Rio (there were no chikungunya clusters found in the west). In contrast, other regions contained many clusters of each disease, though often at different points in time. However, they did not investigate how these differences in clustering might depend on covariates or other factors. Schmidt et al. (2022), using a spatial multinomial regression model, found that cases of chikungunya were more likely in urban areas and Zika cases were less likely in population-dense areas, compared to dengue cases. However, a spatial analysis cannot investigate how space-time factors might have affected differences in the transmission intensity of the diseases, which is of great interest in this example. For instance, laboratory studies have shown that the transmission of Zika may be more sensitive to temperature compared to dengue, which is important as it implies Zika will not be able to spread as effectively into North America and the highlands (Tesla et al., 2018).
Therefore, we apply our MS-ZIARMN model to cases of the three diseases, $K=3$, across the 160 neighborhoods of Rio de Janeiro, $N=160$, for each bi-week between 2015-2016, $T=52$. Our model seems appropriate for this example, as there appear to be long periods of disease absence in both the time series of chikungunya and Zika cases, see Figure 1 for example (overall 65% of Zika cases are equal to 0 and 75% of chikungunya cases are equal to 0).